# Finite Neural Networks as Mixtures of Gaussian Processes: From Provable Error Bounds to Prior Selection

### v1

@@ A.3 Proofs Section 6
${\mathbb{W}}_{\rho}(p_{nn({\mathcal{X}}),k+1},q_{nn({\mathcal{X}}),k+1})$ $\leq\hat{\mathcal{S}}_{\rho,{{w}},k}\left[{\mathcal{L}}_{\sigma}\hat{{\mathbb{W}}}_{\rho,k}+{\mathcal{L}}_{\sigma}\hat{{\mathbb{W}}}_{\rho,{R},k}+\hat{{\mathbb{W}}}_{\rho,{\Delta},k}\right]$
$\coloneqq\hat{{\mathbb{W}}}_{\rho,k+1}$
${\mathbb{W}}_{\rho}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})})\leq\hat{{\mathbb{W}}}_{\rho,K+1}.$
Theorem 17: Uniform Convergence of the Wasserstein Distance from the GMM approximation of a SNN
The proof of Theorem 17 relies on the fact that the $\rho$-Wasserstein distance resulting from the signature operation performed on a Gaussian mixture distribution uniformly converges as the signature size increases, as shown in the following Lemma, which is an extension of Lemma 6.1 Graf and Luschgy (2007), and Corollary.
Lemma 29
(Uniform Convergence in Wasserstein Distance for of a Grid of Signatures) Let $p\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{n})$, then, for $\epsilon>0$ there exists signature locations ${\mathcal{C}}=\{{{c}}_{1},\ldots,{{c}}_{N}\}\subset{\mathbb{R}}^{d}$ that form an $n-$D hyper-cubic lattice in ${\mathbb{R}}^{n}$, i.e., with corresponding regions satisfying Eqn. (10) for some matrix ${{T}}\in{\mathbb{R}}^{d\times d}$, such that
${\mathbb{W}}_{\rho}^{\rho}({\Delta}_{\mathcal{C}}\#p,p)\leq\epsilon.$
Proof  Consider a hyper-cube $\mathcal{A}\subset{\mathbb{R}}^{d}$ such that
[…]
which is guaranteed to exist since $p\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{n})$. Now consider a uniform partition of ${\mathcal{A}}$ in $N$ regions $\{{\mathcal{A}}_{1},...,{\mathcal{A}}_{N}\}$ of diameter $\bar{\epsilon}\leq\frac{1}{\sqrt{n}}\left(\frac{\epsilon}{2}\right)^{1/\rho}$ and for each region ${\mathcal{A}}_{k}$ consider $c_{k}\in{\mathcal{A}}_{k}$ the centroid of ${\mathcal{A}}_{k}$. We let ${\mathcal{C}}=\{{{c}}_{1},{{c}}_{2},\ldots,{{c}}_{N}\}$. Then, the following holds:
${\mathbb{W}}_{\rho}^{\rho}({\Delta}_{\mathcal{C}}\#p,p)$ $=\inf_{\gamma\in\Gamma({\Delta}_{\mathcal{C}}\#p,p)}{\mathbb{E}}_{(\tilde{{{z}}},{{z}})\sim\gamma}[\|\tilde{{{z}}}-{{z}}\|^{\rho}]$
$\leq{\mathbb{E}}_{{{z}}\sim p}[\|{{z}}-\sum_{k=1}^{N}{{c}}_{k}{\mathbf{1}}_{{\mathcal{A}}_{k}}\|^{\rho}]$
[…]
$\leq\epsilon.$
In the last step, we use that the smallest ball that encloses hyper-cube ${\mathcal{A}}_{k}$ has radius $\bar{\epsilon}\sqrt{d}=(\epsilon/2)^{1/\rho}$, so that every point in ${\mathcal{A}}_{k}$ is $(\epsilon/2)$ close to ${{c}}_{k}$ in terms of Euclidean distance for every $k$. Finally, observe that ${\mathcal{C}}$ naturally satisfies constrain Eqn. (10) so that the Lemma is proven by the existence of ${\mathcal{C}}$.
Corollary 30 (of Lemma 29 for Mixture Distributions)
Let $\sum_{i=1}^{M}{{{\pi}}^{(i)}}p_{i}\in{\mathcal{P}}_{2}({\mathbb{R}}^{d})$. Then, for $\epsilon>0$, there exists a set of signature locations ${\mathcal{C}}=\{{\mathcal{C}}_{1},\ldots,{\mathcal{C}}_{M}\}\subset{\mathbb{R}}^{d}$ where every ${\mathcal{C}}_{i}$ satisfies constraint Eqn. (10) for some matrix ${{T}}_{i}\in{\mathbb{R}}^{d\times d}$, such that
${\mathbb{W}}_{2}^{2}\left(\sum_{i=1}^{M}{{{\pi}}^{(i)}}p_{i},\sum_{i=1}^{M}{{{\pi}}^{(i)}}{\Delta}_{{\mathcal{C}}_{i}}\#p_{i}\right)\leq\epsilon.$ (34)
Proof  According to Proposition 24:
${\mathbb{W}}_{2}^{2}\left(\sum_{i=1}^{M}{{{\pi}}^{(i)}}p_{i},\sum_{i=1}^{M}{{{\pi}}^{(i)}}{\Delta}_{{\mathcal{C}}_{i}}\#p_{i}\right)\leq\sum_{i=1}^{M}{{{\pi}}^{(i)}}\left({\mathbb{W}}_{2}^{2}\left(p_{i},{\Delta}_{{\mathcal{C}}_{i}}\#p_{i}\right)\right)$
Consequently, if for all $i\in\{1,\ldots,M\}$ we have that ${\mathbb{W}}_{2}^{2}\left(p_{i},{\Delta}_{{\mathcal{C}}_{i}}\#p_{i}\right)\leq\epsilon/{{{\pi}}^{(i)}}$, then we can conclude Eqn. (34). By Lemma 29, there always exists a ${\mathcal{C}}_{i}$ that satisfies constraint Eqn. (10) such that ${\mathbb{W}}_{2}^{2}\left(p_{i},{\Delta}_{{\mathcal{C}}_{i}}\#p_{i}\right)\leq\epsilon/{{{\pi}}^{(i)}}$, which concludes the proof.
With the results of Lemma 29 and Corollary 30 established, we proceed to prove Theorem 17.
Proof  We prove the Theorem for the $\rho-$Wasserstein with $\rho\in 1,2$. If we do not compress the Gaussian Mixture approximates, i.e., ${R}_{M}(q_{nn({\mathcal{X}}),k})=q_{nn({\mathcal{X}}),k}$ for every $k$, from Proposition 16:
$\hat{{\mathbb{W}}}_{\rho}$ $\leq\mathcal{S}_{\rho,{{{w}}},K}\left[{\mathcal{L}}_{\sigma}\hat{{\mathbb{W}}}_{\rho,K}+\hat{{\mathbb{W}}}_{\rho,{\Delta},K}\right]$
$\leq{\mathcal{L}}_{\sigma}\mathcal{S}_{\rho,{{{w}}},K}\mathcal{S}_{\rho,{{{w}}},K-1}\left[{\mathcal{L}}_{\sigma}\hat{{\mathbb{W}}}_{\rho,K-1}+\hat{{\mathbb{W}}}_{\rho,{\Delta},K-1}\right]+\mathcal{S}_{\rho,{{{w}}},K}{\mathbb{W}}_{\rho,{\Delta},K}$
[…]
$\leq{\mathcal{L}}_{\sigma}^{K}\prod_{l=0}^{k}\mathcal{S}_{\rho,{{{w}}},K-l}\left[{\mathcal{L}}_{\sigma}\hat{{\mathbb{W}}}_{\rho,0}+\hat{{\mathbb{W}}}_{\rho,{\Delta},K-1}\right]+\sum_{k=0}^{K-1}{\mathcal{L}}_{\sigma}^{k}{\mathbb{W}}_{\rho,{\Delta},k}\prod_{l=0}^{k}\mathcal{S}_{\rho,{{{w}}},K-l}$
$=\sum_{k=0}^{K}{\mathcal{L}}_{\sigma}^{k}{\mathbb{W}}_{\rho,{\Delta},k}\prod_{l=0}^{k}\mathcal{S}_{\rho,{{{w}}},K-l},$
Recall that we defined $\mathcal{S}_{\rho,{{w}},k}=D^{\frac{1}{\rho}}{\mathbb{E}}_{{{{W}}}_{k}\sim p_{w_{k}}}[\|{{{W}}}_{k}\|^{\rho}]^{\frac{1}{\rho}}$ (Eqn. (23)), so that, since it is assumed that $p_{{{w}}_{k}}\in{\mathcal{P}}_{\rho}({\mathbb{R}}^{n_{k}\times n_{n_{k-1}}})$, we have that $\mathcal{S}_{\rho,{{w}},k}\in(0,\infty)$ for every $k$. Consequently, it is enough to prove that for any $\epsilon_{\Delta}>0$ there exists a signature location ${\mathcal{C}}_{k}\subset{\mathbb{R}}^{n_{k}}$ of size $N_{k}$ such that for all $k\in\{0,\ldots,K\}$ it holds that $\hat{{\mathbb{W}}}_{\rho,{\Delta},k}<\epsilon_{\Delta}.$ In fact, we can then conclude by taking
$\epsilon_{\Delta}<\frac{\epsilon}{\sum_{k=0}^{K}{\mathcal{L}}_{\sigma}^{k}\prod_{l=0}^{k}\mathcal{S}_{\rho,{{{w}}},K-l}}.$
Let $\hat{{\mathbb{W}}}_{\rho,{\Delta},k}$ be computed as in Eqn. (12) using Corollary 10 and Corollary 27. Then, according to Corollary 30, there exists a set of signature locations ${\mathcal{C}}_{k}$ of finite size, such that $\hat{{\mathbb{W}}}_{\rho,{\Delta},k}<\epsilon_{\Delta}$ for all $k\in\{0,\ldots,k\}$.
A.4 Proofs Section 7
Corollary 21: Wasserstein Distance between any GMM and a SNN
Proof  Since ${\mathbb{W}}_{2}$ satisfies the triangle inequality, we have:
${\mathbb{W}}_{2}(p_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})})\leq{\mathbb{W}}_{2}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})})+{\mathbb{W}}_{2}(q_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}),$
${\mathbb{W}}_{2}(q_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})})\leq{\mathbb{MW}}_{2}(q_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}).$
Combining these two inequalities, we obtain:
${\mathbb{W}}_{2}(p_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})})\leq{\mathbb{W}}_{2}(p_{nn({\mathcal{X}})},q_{nn({\mathcal{X}})})+{\mathbb{MW}}_{2}(q_{nn({\mathcal{X}})},q_{gmm({\mathcal{X}})}),$
which completes the proof.
Proposition 19: Wasserstein-Optimal Grid for Signatures of Multivariate Gaussians
Proof  According to Corollary 10:
${\mathbb{W}}_{2}({\Delta}_{\mathrm{Post}(\bar{{\mathcal{C}}^{1}}\times\ldots\times\bar{{\mathcal{C}}}^{n},{{T}}^{-1})+\{{m}\}}\#\mathcal{N}({m},\Sigma),\mathcal{N}({m},\Sigma))=\sum_{j=1}^{n}{{{\lambda}}^{(j)}}{\mathbb{W}}^{2}_{2}({\Delta}_{\bar{{\mathcal{C}}}^{j}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)).$
$\underset{\begin{subarray}{c}\{\bar{{\mathcal{C}}}^{1},\ldots,\bar{{\mathcal{C}}}^{n}\}\subset{\mathbb{R}},\\ \sum_{i=1}^{n}|\bar{{\mathcal{C}}}^{j}|=N\end{subarray}}{\mathrm{argmin}}{\mathbb{W}}_{2}({\Delta}_{\mathrm{Post}(\bar{{\mathcal{C}}^{1}}\times\ldots\times\bar{{\mathcal{C}}}^{n},{{T}}^{-1})+\{{m}\}}\#\mathcal{N}({m},\Sigma),\mathcal{N}({m},\Sigma))$
$\qquad=\underset{\begin{subarray}{c}\{\bar{{\mathcal{C}}}^{1},\ldots,\bar{{\mathcal{C}}}^{n}\}\subset{\mathbb{R}},\\ \sum_{i=1}^{n}|\bar{{\mathcal{C}}}^{j}|=N\end{subarray}}{\mathrm{argmin}}\sum_{j=1}^{n}{{{\lambda}}^{(j)}}{\mathbb{W}}^{2}_{2}({\Delta}_{\bar{{\mathcal{C}}}^{j}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)),$
$\qquad=\underset{\begin{subarray}{c}\{\bar{{\mathcal{C}}}^{1},\ldots,\bar{{\mathcal{C}}}^{n}\}\subset{\mathbb{R}},\\ \bar{{\mathcal{C}}}^{j}=N_{j}^{*},\forall j\in\{1,\ldots,n\}\end{subarray}}{\mathrm{argmin}}\sum_{j=1}^{n}{{{\lambda}}^{(j)}}{\mathbb{W}}^{2}_{2}({\Delta}_{\bar{{\mathcal{C}}}^{j}}\#\mathcal{N}(0,1),\mathcal{N}(0,1)),$
=$\qquad=\Big{\{}\underset{\bar{{\mathcal{C}}}^{j}\subset{\mathbb{R}},\bar{{\mathcal{C}}}^{j}=N_{j}^{*}}{\mathrm{argmin}}{\mathbb{W}}^{2}_{2}({\Delta}_{\bar{{\mathcal{C}}}^{j}}\#\mathcal{N}(0,1),\mathcal{N}(0,1))\Big{\}}_{j=1}^{n},$
where in step 2 we used the optimal grid-configuration $\{N^{*}_{1},\ldots,N^{*}_{n}\}$ as defined in Eqn.(27), and in step 3, we use the independence of objective terms across dimensions.
Appendix B Compression for Mixtures of Multivariate Bernoulli Distributions
We present an alternative to the procedure in Algorithm 3 to perform the compression operation in the case of mixtures of multivariate Bernoulli distributions. We first show that mixtures of Bernoulli distributions naturally arise in dropout networks and that efficient compression operations are required to formally analyze the output distribution of these networks. We then introduce an alternative compression procedure that includes bounds on the resulting Wasserstein distance.
Let $\phi_{{b}}:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}^{n}$ with binary vector ${{b}}\in\{0,1\}^{n}$ return a vector with the elements in ${{b}}$ masked, i.e., $\phi_{{b}}({{z}})={{b}}\otimes{{z}}$, where $\otimes$ denotes the element-wise multiplication operation. For the dropout operation, we place multi-variate Bernoulli distribution over ${{b}}$ that we denote as $d_{{b}}$. The distribution $d_{{b}}$ can be written:
$d_{{b}}=\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{b}},$
which is parameterized by the success probability, i.e., dropout rate, $\theta\in[0,1]$. For a discrete input distribution $d_{in}=\sum_{i=1}^{N}{{{\pi}}^{(i)}}{{c}}_{i}$, which can represent a (set of) fixed input points of the SNN, the distribution at an intermediate layer of a dropout network, or the signature approximation at an intermediate layer of a VI network, the output distribution of a dropout layer takes the form of a mixture of multivariate Bernoulli distributions:
$d_{out}={\mathbb{E}}_{{{z}}\sim d_{in}}[\phi({{z}})\#d_{{b}}]=\sum_{i=1}^{N}{{{\pi}}^{(i)}}\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{{b}}\otimes{{c}}_{i}},$ (35)
Note that $d_{out}$ is a discrete distribution with a support of size $N\cdot 2^{n}$, i.e., $d_{out}\in{\mathcal{D}}_{N\cdot 2^{n}}({\mathbb{R}}^{n})$. As such, performing formal analysis on $d_{out}$ quickly becomes computationally intractable for reasonably large $n$. Hence, we aim to construct a tractable discrete approximation $\bar{d}_{out}\in{\mathcal{D}}_{M}({\mathbb{R}}^{n})$ of size $M$ of $d_{out}$ with error bounds. For this, we can employ Algorithm 3, which clusters the elements of $d_{out}$ into $M$ clusters using M-means clustering and replaces each cluster with a single element using moment matching. The algorithm then solves a discrete optimal transport problem to compute the resulting 2-Wasserstein distance. This procedure can be computationally demanding for mixtures with a large number of components. Instead, we can use the structure of the mixture of Bernoulli distributions to efficiently cluster the mixtures and obtain a closed-form upper bound on the 2-Wasserstein distance, as we show below.
To compress $d_{out}$ to a discrete distribution with $M$ elements, we first collect the indices of the $\log_{2}(M)$ neurons with the largest absolute weight value in set ${\mathcal{I}}$ using a simple sorting operation. Then, we take
$\bar{d}_{out}=\sum_{i\in{\mathcal{I}}^{c}}{{{\pi}}^{(i)}}\delta_{{{c}}_{i}}+\sum_{i\in{\mathcal{I}}}{{{\pi}}^{(i)}}\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{{b}}\otimes{{c}}_{i}}$ (36)
where ${\mathcal{I}}^{c}=\{1,\ldots,N\}\setminus{\mathcal{I}}$. Note that $\bar{d}_{out}$ is equivalent to a mixture of multivariate Bernoulli distributions with outcomes mapped to ${{\bar{0}}}$ and ${{c}}_{i}$ and with probability parameter $1$ for the dimensions in ${\mathcal{I}}^{c}$ and $\theta$ for the dimensions in ${\mathcal{I}}$. The following proposition provides a tight closed-form upper-bound on the 2-Wasserstein distance between $d_{out}$ and $\bar{d}_{out}$.
Proposition 31
For multivariate Bernoulli distributions $d_{out}\in{\mathcal{D}}_{N\cdot 2^{n}}({\mathbb{R}}^{n})$ defined as in Eqn. (35) and $\bar{d}_{out}\in{\mathcal{D}}_{M}({\mathbb{R}}^{n})$ as defined in Eqn. (36), we have that
${\mathbb{W}}_{2}(d_{out},\bar{d}_{out})\leq\sum_{i\in{\mathcal{I}}^{c}}^{N}(1-\theta)\|{{c}}_{i}\|^{2}$
Proof  According to Proposition 24:
${\mathbb{W}}_{2}^{2}(d_{out},\bar{d}_{out})$
$\qquad\leq\sum_{i\in{\mathcal{I}}^{c}}{{{\pi}}^{(i)}}{\mathbb{W}}_{2}^{2}\big{(}\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{{b}}\otimes{{c}}_{i}},\delta_{{{c}}_{i}}\big{)}$
$\qquad+\sum_{i={\mathcal{I}}}{{{\pi}}^{(i)}}{\mathbb{W}}_{2}^{2}\big{(}\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{{b}}\otimes{{c}}_{i}},\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{{b}}\otimes{{c}}_{i}}\big{)}$
Since the Wasserstein distance between a distribution and itself is zero, we obtain:
${\mathbb{W}}_{2}^{2}(d_{out},\bar{d}_{out})\leq\sum_{i\in{\mathcal{I}}^{c}}{{{\pi}}^{(i)}}{\mathbb{W}}_{2}^{2}\big{(}\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{{b}}\otimes{{c}}_{i}},\delta_{{{c}}_{i}}\big{)}$
The trivial optimal transport plan for each $i$ is to move all probability mass from $\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{{b}}\otimes{{c}}_{i}}$ to ${{c}}_{i}$. Therefore, for each $i$, we have that:
${\mathbb{W}}_{2}^{2}\big{(}\sum_{{{b}}\in\{0,1\}^{n}}\theta^{\sum_{l=1}^{n}{{{b}}^{(l)}}}(1-\theta)^{n-\sum_{l=1}^{n}{{{b}}^{(l)}}}\delta_{{{b}}\otimes{{c}}_{i}},\delta_{{{c}}_{i}}\big{)}=(1-\theta)\|{{c}}_{i}\|$
which completes the proof.
Appendix C Results on Signatures of Gaussian Mixture Distributions
(a) Empirical estimates and formal bounds on $\overline{W}_{2}$ from grid-constrained signatures.
(b) Empirical estimates of $\overline{W}_{2}$ from grid-constrained and unconstrained signature locations.
Figure 9: The relative 2-Wasserstein distance between a 2D Gaussian mixture distribution with $M$ components and the signature with $N$ grid-constrained locations obtained via Algorithm 2 (solid and dotted lines), and the signature with $N$ unconstrained optimal locations (dash-dotted line). The unconstrained optimal locations are obtained by performing $N$-Means clustering on $31e$ 3 MC samples from the GMM. The solid lines represent the formal bounds on the relative 2-Wasserstein distance provided by Algorithm 2, while the dotted and dash-dotted lines show empirical approximations.
In this section, we experimentally evaluate the effectiveness of Algorithm 2 in constructing signatures on Gaussian mixtures. Specifically, we examine the conservatism of the formal bound on the 2-Wasserstein distance resulting from the signature operation provided by Algorithm 2, and analyze the restrictiveness of having the signature locations of each component of the mixture on a grid as in Algorithm 2.
In Figure 9(a), we analyze how the approximation error from the signature of a Gaussian mixture with $M$ components, obtained according to Algorithm 2, changes with increasing signature size. The plots show that as the signature size increases, the approximation error, measured as the relative 2-Wasserstein distance $\overline{W}_{2}$, decreases uniformly. Furthermore, in line with Corollary 10, the formal upper bound on $\overline{W}_{2}$ is exact for Gaussians ($M=1$). For Gaussian mixtures ($M>1$), the conservatism introduced by the formal bounds grows approximately linearly with increasing $M$. This is because, in Algorithm 1, we bound the Wasserstein distance resulting from the signature operation on the GMMs by the weighted sum of the 2-Wasserstein distance between each Gaussian component in the mixture and their signatures (see line 7 in Algorithm 2 and Eqn. (12)).
Recall from Subsection 5.2 that to obtain a closed-form bound on the approximation error in Algorithm 2, we place the signature locations of each component of the mixture on a grid in the transformed space induced by the element’s covariance matrix, as illustrated in Figure 4. Specifically, following Proposition 19, we choose the grids such that for a given signature size $N$, the 2-Wasserstein distance from the signature operation is minimized. In Figure 9(b), we estimate the optimality gap if we could instead place the signatures freely to minimize the 2-Wasserstein distance. Perhaps surprisingly, the plots show that placing the signatures freely only slightly reduces the 2-Wasserstein distance. Thus, although the optimal signature locations do not adhere to a grid structure (in line with literature (Graf and Luschgy, 2007)), the optimal grid approximates the optimal locations closely.

### v2

@@ Appendix E Compression for Dropout-Induced Bernoulli Mixtures
So, ${\mathbb{W}}_{2}^{2}\left(p_{i},\delta_{{{c}}_{i}}\right)=(1-\theta)\|{{c}}_{i}\|^{2}$, which completes the proof.
Appendix F Results on Signatures of Gaussian Mixture Distributions
Here, we experimentally evaluate the effectiveness of Algorithm 2 in constructing signatures on Gaussian mixtures. Specifically, we examine the conservatism of the formal bound on the 2-Wasserstein distance resulting from the signature operation provided by Algorithm 2, and analyze the restrictiveness of having the signature locations of each component of the mixture on a grid as in Algorithm 2.
Figure 11: Empirical estimates (dashed lines) and formal bounds (solid lines) on ${\mathbb{W}}_{2}$ between a 2D Gaussian mixture distribution with $M$ components and the signature with NN grid-constrained locations obtained via Algorithm 2. Formal bounds are computed with Algorithm 2. Empirical estimates are computed from $10^{4}$ samples from each distribution and computing the 2‑Wasserstein distance between the resulting empirical distributions.
In Figure 11, we analyze how the approximation error from the signature of a Gaussian mixture with $M$ components, obtained according to Algorithm 2, changes with increasing signature size. The plots show that as the signature size increases, the approximation error, measured as the 2-Wasserstein distance, decreases uniformly. Furthermore, in line with Corollary 11, the formal upper bound on ${\mathbb{W}}_{2}$ is exact for Gaussians ($M=1$). For Gaussian mixtures ($M>1$), the conservatism introduced by the formal bounds grows approximately linearly with increasing $M$. This is because, in Algorithm 1, we bound the Wasserstein distance resulting from the signature operation on the GMMs by the weighted sum of the 2-Wasserstein distance between each Gaussian component in the mixture and their signatures (see line 7 in Algorithm 2 and Eqn (14)).
Recall from Subsection 5.2 that to obtain a closed-form bound on the approximation error in Algorithm 2, we place the signature locations of each component of the mixture on a grid in the transformed space induced by the element’s covariance matrix, as illustrated in Figure 4. Specifically, following Proposition 23, we choose the grids such that for a given signature size $N$, the 2-Wasserstein distance from the signature operation is minimized. In Figure 5(b), we estimate the optimality gap if we could instead place the signatures freely to minimize the 2-Wasserstein distance. Perhaps surprisingly, the plots show that placing the signatures freely only slightly reduces the 2-Wasserstein distance. Thus, although optimal signature locations are not grid-aligned (in line with literature (Graf and Luschgy, 2000)), for the Gaussian mixtures with distinct modes considered here the optimally chosen grid closely approximates the unconstrained optimal locations, indicating that component-wise grids can be near-optimal for such mixtures.